# Learning holographic QCD with unflavoured meson spectra

### v1

@@ Abstract
We develop a neural network framework to predict the five-dimensional background geometry, dilaton potential, and chiral symmetry breaking scalar potential of holographic QCD from unflavored meson mass spectra. The model was trained in a discretized form of the Schrödinger-like equation, which resembles a linear moose in “deconstructed” 5 dimensions with Dirichlet boundary conditions, in contrast to the AdS/DL with “emergent” space-time. Using the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ unflavored mesons and their excitations as training data, the model learns confining effective potentials and computes a dilaton profile that satisfies the null energy condition. The network predicts the IR behavior of dilaton to be in-between linear and quadratic forms. Moreover, the symmetry-breaking bulk potential of the scalar field, $V(X)=k_{1}X^{3}+k_{2}X^{4}$, was computed, and the parameters $k_{1}$ and $k_{2}$ predicted to be $\sim-\ 8$ and $\sim 17$ respectively. The deep-learned parameters, metric, and the dilaton profile were then used to predict the pion mass and its spectrum with good accuracy. A Python code, along with the trained models, is provided to facilitate further studies [^3] Available at Github, https://github.com/rp-winter/NN-AdS-QCD.
@@ 1 Introduction
The AdS/CFT correspondence [1, 2, 3] has provided a powerful framework to study strongly coupled gauge theories using classical gravitational theories in higher-dimensional spacetime. One of the most prominent applications of this correspondence is in the study of Quantum Chromodynamics (QCD) through holographic models. The original proposal by Sakai and Sugimoto [4, 5] with $D8/D4/\overline{D8}$ configuration of the branes to achieve chiral symmetry and its spontaneous breaking, despite its success, could not produced massive pions. This geometry was deformed to $D8/D4/\overline{D8}/D6$ [6, 7], to mimic the techni-colour QCD like sector on the $D6$ brane, which successfully provided explicit chiral symmetry breaking through a bare quark mass term arising from the asymptotic distance between the $D4$ and $D6$ branes. While such techni-colour models, for chiral symmetry breaking, have rich phenomenology and were studied in the context of raising the axion mass [8, 9], with the discovery of Higgs, remains disfavoured.
@@
The bottom-up AdS/QCD framework [10, 11, 12, 13, 14, 15, 16], in contrast to the string-theoretic approach, remains successful in explaining phenomenological signatures, including the mass spectra of mesons and their decay constants. Unlike full string-theoretic duality, parameters and interactions are chosen to reproduce low-energy QCD observables phenomenologically. These models aim to capture the essential features of QCD, such as confinement and chiral symmetry breaking, by constructing a dual gravitational theory in a 5-dimensional Anti-de Sitter (AdS) space related to the 4-dimensional gauge theory through an effective dictionary. Although the exact gravity dual of QCD and the chiral symmetry breaking dynamics remain unknown, phenomenological models successfully reproduce various hadronic properties, such as the meson mass spectrum [12, 17, 18, 19], which are identified with the Kaluza-Klein (KK) modes of the bulk fields. The dynamics of these bulk fields are governed by Sturm-Liouville-type differential equations, and their eigenvalues correspond to these masses in the dual theory. These equations can often be transformed into Schrödinger-like equations [12], allowing the application of quantum mechanical techniques [20].
@@
There are several proposals for the functional forms of the background fields (such as dilaton profile, warp factor, and scalar vacuum expectation value), each characterized by a few free parameters, which are then fitted to match the hadronic spectra. For example, the Gherghetta-Kaputsa-Kelley (GKK) model [17] takes the standard AdS warp factor A(z)=log(z)A(z)=-log(z), and assumes an ansatz for the bulk scalar vacuum expectation value (VEV) field to reconstruct the dilaton profile. In many improved holographic QCD models, with 5D graviton–dilaton (or graviton–dilaton–scalar) action [21, 22, 23], solving the corresponding Einstein-dilaton equations generate a self-consistent warp factor and dilaton profile, instead of an ad hoc selection.
@@
Although these models can reproduce the low-lying hadronic spectra with reasonable accuracy, they are often limited by the assumptions made about these background field profiles. Neural networks, being a universal function approximator [24], can be used to approximate these background fields without making any assumptions about their functional forms. There has been an increased interest in the subject with the establishment of a correspondence between the AdS/CFT and Deep Learning (DL) [25], where the deep layers of a deep Boltzmann machine, from the viewpoint of a 4-dimensional QFT at the visible layer, are seen as an “emergent” bulk space coordinate. The trained weights become the metric of the extra-dimensional space-time. Further, the matching of the DL with holographic QCD [26, 27], leverages this correspondence to model the background geometry by fitting the $\rho$ and $a_{2}$ mesons using neural networks. This framework was then improved by including a dilaton potential to predict the rho meson spectrum [28].
@@
$\rho$ 1 $0.775\pm 0.023$
2 $1.465\pm 0.025$
3 $1.57\pm 0.07$
4 $1.72\pm 0.02$
5 $1.88\pm 0.01$
6 $2.044\pm 0.031$
$a_{1}$ 1 $1.23\pm 0.04$
2 $1.655\pm 0.016$
$a_{2}$ 1 $1.318\pm 0.006$
2 $1.706\pm 0.014$
$f_{0}$ 1 $0.513\pm 0.032$
2 $0.99\pm 0.02$
3 $1.40\pm 0.04$
4 $1.522\pm 0.025$
5 $1.733\pm 0.008$
6 $1.804\pm 0.016$
7 $1.982\pm 0.54$
8 $2.187\pm 0.014$
9 $2.312\pm 0.007$
10 $2.470\pm 0.006$
Table 1: Experimental meson masses[47] used for training the neural network.
@@
In this article, we train a neural network model to learn the background geometry of the AdS/QCD model, the chiral symmetry breaking potential, dilton profile and the effective potentials of the mesons by fitting the mass spectrum of the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons and all their excitations, given in Table 1. The mass eigenvalues are obtained by solving the Schrödinger-like equation for the bulk modes using a finite-difference discretization, where the spatial coordinate is divided into N points and the second derivative is approximated by a central difference scheme. This leads to a matrix eigenvalue problem, which can often be cast into a real symmetric form and solved using numerical eigenvalue solvers. By minimizing the difference between the predicted and experimental meson masses using gradient descent, we can optimize the neural networks to find the background geometry. The Python code can be found at [29], which can be easily modified to train further models. In contrast with the AdS/DL framework, the neural network model we construct here closely follows the “deconstruction” [30, 31, 32, 33, 34, 35, 36, 37] of a warped 5-dimensional theory with Dirichlet boundary conditions [38].
@@
Our results show good agreement with the experimental values for the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ meson masses used in training. To test the generalization capability of the trained model, we use the same trained model to predict the excited-state masses of the $\pi$ mesons, which are in reasonable agreement with the experimental values, demonstrating the effectiveness of our approach in capturing the underlying physics of AdS/QCD. In contrast with the GKK model [17], the dilaton profile we compute remains positive throughout the 5-dimensional space, which is consistent with the null energy condition [39, 16, 40]. The dilaton profile we obtain lies in-between the linear and quadratic functions. This finding is crucial because, though string theoretical models predict a linear dilaton profile, it would result in a continuum spectrum for the glueballs [41]. The dilaton wave profile, along with the scalar vevs, provides a deep enough potential for the pion wavefunction to support a $\sim 135$ MeV light state.
@@
In Section.2, we briefly introduce the bottom-up AdS/QCD, including the scalar, pseudo-scalar, vector, axial-vector, and higher spin mesons. In Section.3, we discuss the methodology and define the Loss function for the neural network. Subsequently, in Section.4 we give the numerical results and predictions, and we conclude in Section 5.
@@ 2 Review of AdS/QCD
where, $z$ is defined as a dimensionless parameter related the the physical scale up to the radius of the AdS space ($L$) and the Minkowski metric is chosen to be $\eta_{\mu\nu}=\text{diag}(-1,+1,+1,+1)$.
@@
where, $m_{X}^{2}=-3$, $g_{5}^{2}=12\pi^{2}/N_{c}$ ($N_{c}$ being the number of colors), $D^{M}X=\partial^{M}X-iA_{L}^{M}X+iXA_{R}^{M}$ is the covariant derivative, and $F_{L,R}^{MN}=\partial^{M}A_{L,R}^{N}-\partial^{N}A_{L,R}^{M}-i\left[A_{L,R}^{M},A_{L,R}^{N}\right]$ are the field tensors. The scalar potential is taken to be $V(X)=\frac{4}{3}k_{1}X^{3}+2k_{2}X^{4}$ [12].
@@
Varying the action, the dependence of VEV on the 5 dimensional space is given by,
@@
where, $V_{k}(v(z))=\frac{\partial V(X)}{\partial X}\Big|_{X=v(z)}=k_{1}v(z)^{2}+k_{2}v(z)^{3}$.
@@ 2.1.1 Scalar and Pseudo-Scalar Mesons
$\partial_{z}\left(e^{-B_{2}(z)}\partial_{z}S_{n}(z)\right)-e^{-B_{2}(z)}e^{2A(z)}\left(m_{X}^{2}-(2k_{1}v(z)+3k_{2}v^{2}(z))\right)S_{n}(z)=-e^{-B_{2}(z)}m^{2}_{Sn}S_{n}(z),$ (8)
where, $B_{2}(z)=\phi(z)-3A(z)$, and $m_{Sn}$ are the discrete eigenvalues of the Sturm-Liouville problem. These discrete values of $m_{Sn}^{2}$ correspond to the mass of the scalar mesons. This equation can be transformed into a Schrödinger equation, using the substitution $S_{n}(z)=e^{-B_{2}(z)/2}\psi_{n}(z)$, giving,
$\partial_{z}^{2}\psi_{n}(z)+V_{f_{0}}(z)\psi_{n}(z)=m^{2}_{Sn}\psi_{n}(z)\ ,$ (9)
@@
Upon investigating the UV behavior of the potential $V_{f_{0}}$, we observe that $V_{f_{0}}(z\rightarrow 0)\rightarrow+\infty$, which ensures that the wave-function $\psi_{n}(z)$ goes to zero at the UV boundary, along with the normalizability condition $\psi_{n}(z\rightarrow\infty)\rightarrow 0$.
@@
$\frac{e^{B_{2}(z)}}{v^{2}(z)}\partial_{z}\left(v^{2}(z)e^{-B_{2}(z)}\partial_{z}\pi_{n}\right)+m^{2}_{\pi n}(\pi_{n}-\xi_{n})=0\ ,$ (11)
here, we work in the gauge $A_{5}=0$. The longitudinal component of the axial vector field, $A_{\mu}=(A_{L\mu}-A_{R\mu})/2$, is denoted by $\xi(x,z)=\sum_{n}\Xi_{n}(x)\xi_{n}(z)$, where $n$ denotes the KK mode.
@@
where $B_{1}(z)=\phi(z)-A(z)$. To solve these coupled differential equations, it is possible to combine Eq.11 and Eq.12, into a Schrödinger-like equation [19], but doing so eliminates information from the mass spectrum [43, 44]. Specifically, we will only be able to get the mass eigenvalues corresponding to the excited states, and lose information of the ground state mass. We therefore choose to solve these two coupled second-order differential equations directly. Simplifying these equations, we get,
$\begin{split}\partial_{z}^{2}\pi_{n}+\omega(z)\partial_{z}\pi_{n}+m_{\pi n}^{2}(\pi_{n}-\xi_{n})&=0,\\ \partial_{z}^{2}\xi_{n}-(\partial_{z}B_{1}(z))\partial_{z}\xi_{n}+C(z)(\pi_{n}-\xi_{n})&=0,\end{split}$ (13)
where $\omega(z)=2\partial_{z}v(z)/v(z)-\partial_{z}B_{2}(z)$ and $C(z)=g_{5}^{2}v^{2}(z)e^{2A(z)}$. These two equations can be solved with the boundary conditions $\pi_{n}=0$ and $\xi_{n}=0$ at $z\rightarrow 0$ and $z\rightarrow\infty$ to obtain the mass spectrum of the pseudo-scalar mesons.

### v2

@@ Abstract
We develop a data-driven neural network framework to reconstruct the five-dimensional background geometry, the dilaton potential, and the chiral-symmetry-breaking scalar potential of holographic QCD from hadron mass spectra. Framed as an inverse problem, the model is trained using a discretized form of the Schrödinger-like equation, which resembles a linear moose in “deconstructed” 5 dimensions with Dirichlet boundary conditions, in contrast to the AdS/DL with “emergent” space-time. Using the masses of the unflavored mesons $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ and their excitations as training data, the model learns confining effective potentials and computes a dilaton profile that satisfies the null energy condition. The network predicts that the dilaton’s IR behavior will be much steeper than its quadratic form. Moreover, the symmetry-breaking bulk potential of the scalar field, $V(X)=k_{1}X^{3}+k_{2}X^{4}$, was computed, and the parameters $k_{1}$ and $k_{2}$ predicted to be $\sim-4$ and $\sim 9$ respectively. The deep-learned parameters, metric, and the dilaton profile were then used to predict the pion mass and its spectrum with good accuracy. A Python code, along with the trained models, is provided to facilitate further studies [^3] Available at Github, https://github.com/rp-winter/NN-AdS-QCD.
@@ 1 Introduction
The AdS/CFT correspondence [51, 28, 65] has provided a powerful framework to study strongly coupled gauge theories using classical gravitational theories in higher-dimensional spacetime. One of the most prominent applications of this correspondence is in the study of Quantum Chromodynamics (QCD) through holographic models. The original proposal by Sakai and Sugimoto [59, 60] with $D8/D4/\overline{D8}$ configuration of the branes to achieve chiral symmetry and its spontaneous breaking, despite its success, could not produce massive pions. This geometry was deformed to $D8/D4/\overline{D8}/D6$ [49, 30], to mimic the techni-colour QCD-like sector on the $D6$ brane, which successfully provided explicit chiral symmetry breaking through a bare quark mass term arising from the asymptotic distance between the $D4$ and $D6$ branes. While such techni-colour models, for chiral symmetry breaking, have rich phenomenology and were studied in the context of raising the axion mass [38, 39], they remain disfavoured with the discovery of the Higgs.
@@
The bottom-up AdS/QCD framework [21, 37, 43, 18, 24, 62, 44], in contrast to the string-theoretic approach, remains successful in explaining phenomenological signatures, including the mass spectra of mesons and their decay constants. Unlike full string-theoretic duality, parameters and interactions are chosen to reproduce low-energy QCD observables phenomenologically. These models aim to capture the essential features of QCD, such as confinement and chiral symmetry breaking, by constructing a dual gravitational theory in 5-dimensional Anti-de Sitter (AdS) space, related to the 4-dimensional gauge theory via an effective dictionary. Although the exact gravity dual of QCD and the chiral symmetry breaking dynamics remain unknown, phenomenological models successfully reproduce various hadronic properties, such as the meson mass spectrum [43, 27, 67, 64], which are identified with the Kaluza-Klein (KK) modes of the bulk fields. The dynamics of these bulk fields are governed by Sturm-Liouville-type differential equations, and their eigenvalues correspond to these masses in the dual theory. These equations can often be transformed into Schrödinger-like equations [43], allowing the application of quantum mechanical techniques [66].
@@
There are several proposals for the functional forms of the background fields (such as the dilaton profile, the warp factor, and the scalar vacuum expectation value), each characterized by a few free parameters, which are then fitted to match the hadronic spectra. For example, the Gherghetta-Kaputsa-Kelley (GKK) model [27] takes the standard AdS warp factor A(z)=log(z)A(z)=-log(z), and assumes an ansatz for the bulk scalar vacuum expectation value (VEV) field to reconstruct the dilaton profile. In many improved holographic QCD models, with 5D graviton–dilaton (or graviton–dilaton–scalar) action [20, 11, 52], solving the corresponding Einstein-dilaton equations generates a self-consistent warp factor and dilaton profile, instead of an ad hoc selection.
@@
Although these models can reproduce the low-lying hadronic spectra with reasonable accuracy, they are often limited by the assumptions made about these background field profiles. Neural networks, as universal function approximator [40], can approximate these background fields without making any assumptions about their functional forms. There has been an increased interest in the subject with the establishment of a correspondence between the AdS/CFT and Deep Learning (DL) [35, 42], where the deep layers of a deep Boltzmann machine, from the viewpoint of a 4-dimensional QFT at the visible layer, are seen as an “emergent” bulk space coordinate. The trained weights serve as the metric for the extra-dimensional space-time. Further, the matching of the DL with holographic QCD [33, 7, 31], leverages this correspondence to model the background geometry by fitting the $\rho$ and $a_{2}$ mesons using neural networks. This framework was then improved by including a dilaton potential to predict the rho meson spectrum [32].
Beyond these, there has been growing interest in the applications of machine learning techniques in Holographic QCD [41, 26, 4]. In particular, machine learning has been employed in the study of glueball spectra [15], heavy-quark potentials [50], heavy quarkonium spectra and their decay constants [68], probing proton structure [48], and using lattice QCD data to learn the QCD phase diagram [13]. These developments collectively demonstrate that holographic QCD can be effectively formulated as an inverse problem, with machine learning providing a flexible framework for exploring the space of admissible background configurations.
@@
$\rho$ 11 $0.775\pm 0.001$ [54]
22 $1.465\pm 0.025$ [54]
33 $1.720\pm 0.020$ [54]
$4^{\prime}$ $1.909\pm 0.030$ [10]
$5^{\prime}$ $2.095\pm 0.004$ [3]
$a_{1}$ 1 $1.230\pm 0.040$ [54]
2 $1.655\pm 0.016$ [54]
$a_{2}$ 1 $1.318\pm 0.001$ [54]
2 $1.706\pm 0.014$ [54]
$f_{0}$ 11 $0.478^{+0.029}_{-0.028}$ [5]
22 $0.990\pm 0.020$ [54]
$3^{\prime}$ $1.472\pm 0.012$ [9]
44 $1.522\pm 0.025$ [54]
55 $1.733^{+0.008}_{-0.007}$ [54]
66 $1.982^{+0.054}_{-0.003}$ [54]
$7^{\prime}$ $2.187\pm 0.014$ [54]
$8^{\prime}$ $2.340\pm 0.02$ [61]
$9^{\prime}$ $2.470^{+0.006}_{-0.007}$ [2]
$\pi$ $1^{\prime}$ $0.135\pm 0.001$ [54]
$2^{\prime}$ $1.300\pm 0.100$ [54]
$3^{\prime}$ $1.81^{+0.009}_{-0.011}$ [54]
Table 1: Experimental masses (Breit-Wigner) of the mesons. The unprimed numbers refer to states whose masses we have used for training, and the primed numbers refer to the states that we predict using the trained model.
@@
In this article, we develop a neural network framework to learn the background geometry of the AdS/QCD model, the chiral symmetry breaking potential, dilaton profile, and the effective potentials of the mesons by fitting the mass spectrum of the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ mesons and all their excitations, given in Table 1. The mass eigenvalues are obtained by solving the Schrödinger-like equation for the bulk modes using a finite-difference discretization, where the spatial coordinate is divided into N points and the second derivative is approximated by a central difference scheme. This leads to a matrix eigenvalue problem, which can often be cast into a real symmetric form and solved using numerical eigenvalue solvers. By minimizing the difference between the predicted and experimental meson masses using gradient descent, we can optimize the neural networks to find the background geometry. The Python code, along with the trained models, is available at [56]. The code can be easily modified to train further models, while the trained networks can be reused to explore additional observables and extensions of the AdS/QCD setup. In contrast with the AdS/DL framework, the neural network model we construct here closely follows the “deconstruction” [8, 16, 1, 58, 23, 63, 19, 22] of a warped 5-dimensional theory with Dirichlet boundary conditions [53].
For training, we have used the data of the mean values of the estimated Breit-Wigner (BW) masses from the PDG Summary Table [54]. The Summary table estimates the $f_{0}(500)$ BW mass to be in the range 0.4 to 0.8 GeV, and the T-matrix pole ($\sqrt{s}$) mass to be in the range 0.40 to 0.55 GeV. To train the model, we have used the central value (0.475 GeV) of the T-Matrix Pole for this state. We scale down the loss associated with this mass to reduce its contribution to the total loss. This allows for greater variation in its mass during training. The estimated BW mass of $f_{0}(1370)$ as given in the Summary Table is in the range 1.2 to 1.5 GeV. We do not include this state in the training data and leave it for prediction by our trained model.
@@
Our results show good agreement with the experimental values for the $\rho$, $a_{1}$, $a_{2}$, and $f_{0}$ meson masses used in training. The mass of $f_{0}(1370)$, which is not well determined experimentally, was predicted to be around $1.49\pm 0.02$ GeV, which is consistent with the experimental range of $1.200\text{ to }1.500$ GeV. We further predict the masses of the higher excited states of the $\rho$ and $f_{0}$ mesons, that were not used in training (not included in the Summary Table), whose results are given in Table 3. To assess the generalization capability of the trained model to other meson sectors, we examine its predictions for the excited-state masses of $\pi$ mesons. While the $\pi$ sector is not included in the training loss function, it is used as an additional constraint during hyperparameter optimization. The resulting predictions are in reasonable agreement with experimental values, indicating that the model captures key features of the underlying AdS/QCD dynamics. In contrast with the GKK model [27], the dilaton profile we compute remains positive throughout the 5-dimensional space, which is consistent with the null energy condition [47, 44, 14]. The dilaton profile we obtain increases more steeply than the quadratic functions. This finding is crucial because, though string-theoretical models predict a linear dilaton profile, such a profile would yield a continuum spectrum for glueballs [29]. The dilaton wave profile, along with the scalar vevs, provides a deep enough potential for the pion wavefunction to support a $\sim 135$ MeV light state.
@@
In Section 2, we briefly introduce the bottom-up AdS/QCD, including the scalar, pseudo-scalar, vector, axial-vector, and higher spin mesons. In Section 3, we discuss the methodology and define the Loss function for the neural network. Subsequently, in Section 4, we present the numerical results and predictions, and conclude in Section 5.
@@ 2 Review of AdS/QCD
where, $z$ is defined as a dimensionless parameter related to the physical scale up to the radius of the AdS space ($L$) and the Minkowski metric is chosen to be $\eta_{\mu\nu}=\text{diag}(-1,+1,+1,+1)$.
@@
where, $m_{X}^{2}=-3$, $g_{5}^{2}=12\pi^{2}/N_{c}$ ($N_{c}$ being the number of colors), $D^{M}X=\partial^{M}X-iA_{L}^{M}X+iXA_{R}^{M}$ is the covariant derivative, and $F_{L,R}^{MN}=\partial^{M}A_{L,R}^{N}-\partial^{N}A_{L,R}^{M}-i\left[A_{L,R}^{M},A_{L,R}^{N}\right]$ are the field tensors. The scalar potential is taken to be $V(X)=\frac{4}{3}k_{1}X^{3}+2k_{2}X^{4}$.
@@
Varying the action, the dependence of VEV on the 5-dimensional space is given by,
@@
where, $V_{k}(v(z))=\frac{\partial}{\partial v(z)}\text{Tr}\left(V\left(\langle X\rangle\right)\right)=k_{1}v(z)^{2}+k_{2}v(z)^{3}$.
@@ 2.1.1 Scalar and Pseudo-Scalar Mesons
$\partial_{z}\left(e^{-B_{2}(z)}\partial_{z}S_{n}(z)\right)-e^{-B_{2}(z)}e^{2A(z)}\left(m_{X}^{2}-(2k_{1}v(z)+3k_{2}v^{2}(z))\right)S_{n}(z)=-e^{-B_{2}(z)}m^{2}_{f_{0},n}S_{n}(z),$ (8)
where, $B_{2}(z)=\phi(z)-3A(z)$, and $m_{f_{0},n}^{2}$ are the discrete eigenvalues of the Sturm-Liouville problem. These discrete values of $m_{f_{0},n}$ correspond to the mass of the scalar mesons. This equation can be transformed into a Schrödinger equation, using the substitution $S_{n}(z)=e^{B_{2}(z)/2}\psi_{f_{0},n}(z)$, giving,
$\partial_{z}^{2}\psi_{f_{0},n}+V_{f_{0}}(z)\psi_{f_{0},n}=m^{2}_{f_{0},n}\psi_{f_{0},n}\ ,$ (9)
@@
Upon investigating the UV behavior of the potential $V_{f_{0}}$, we observe that $V_{f_{0}}(z\rightarrow 0)\rightarrow+\infty$, which ensures that the wave-function $\psi_{f_{0},n}(z)$ goes to zero at the UV boundary, along with the normalizability condition $\psi_{f_{0},n}(z\rightarrow\infty)\rightarrow 0$.
@@
$\frac{e^{B_{2}(z)}}{v^{2}(z)}\partial_{z}\left(v^{2}(z)e^{-B_{2}(z)}\partial_{z}\pi_{n}\right)+m^{2}_{\pi,n}(\pi_{n}-\xi_{n})=0\ .$ (11)
Here, we work in the gauge $A_{5}=0$, where the longitudinal component of the axial vector field, $A_{\mu}=(A_{L\mu}-A_{R\mu})/2=A_{\perp\mu}+\partial_{\mu}\xi(x,z)$, is denoted by $\xi(x,z)=\sum_{n}\Xi_{n}(x)\xi_{n}(z)$, where $n$ denotes the KK mode.
@@
where $B_{1}(z)=\phi(z)-A(z)$. To solve these coupled differential equations, it is possible to combine Eq. 11 and Eq. 12, into a Schrödinger-like equation [64], but doing so eliminates information from the mass spectrum [45, 12]. Specifically, we will only be able to obtain the mass eigenvalues of the excited states, losing information about the ground-state mass. We therefore choose to solve these two coupled second-order differential equations directly. Simplifying these equations, we get,
$\begin{split}\partial_{z}^{2}\pi_{n}+\omega(z)\partial_{z}\pi_{n}+m_{\pi,n}^{2}(\pi_{n}-\xi_{n})&=0,\\ \partial_{z}^{2}\xi_{n}-(\partial_{z}B_{1}(z))\partial_{z}\xi_{n}+C(z)(\pi_{n}-\xi_{n})&=0,\end{split}$ (13)
where $\omega(z)=2\partial_{z}v(z)/v(z)-\partial_{z}B_{2}(z)$ and $C(z)=g_{5}^{2}v^{2}(z)e^{2A(z)}$. These two equations can be solved with the boundary conditions $\pi_{n}=\xi_{n}=0$ at $z\rightarrow 0$, and $\partial_{z}\pi_{n}=\partial_{z}\xi_{n}=0$ at $z\rightarrow\infty$ to obtain the mass spectrum of the pseudo-scalar mesons.